40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 00:07:54 2015, rana, Update, LSC, Lock Loss plot 1107673198.png1107673198.png
    Reply  Wed Feb 11 02:42:05 2015, Jenne, Update, LSC, PRC error signal RF spectra REFL33_25.pdfREFL33_15.pdf
       Reply  Wed Feb 11 03:41:12 2015, Koji, Update, LSC, PRC error signal RF spectra 
          Reply  Wed Feb 11 17:31:11 2015, ericq, Update, LSC, RFPD spectra REFL.pngAS.pngPOP.png2015-02-PDSpectra.zip
             Reply  Wed Feb 11 18:07:42 2015, ericq, Update, LSC, RFPD spectra 
                Reply  Thu Feb 12 01:00:18 2015, rana, Update, LSC, RFPD spectra 
Message ID: 11000     Entry time: Wed Feb 11 03:41:12 2015     In reply to: 10999     Reply to this: 11003
Author: Koji 
Type: Update 
Category: LSC 
Subject: PRC error signal RF spectra 

As the measurements have been done under feedback control, the lower RF peak height does not necessary mean
the lower optical gain although it may be the case this time.

These non-33MHz signals are embarassingly high!
We also need to check how these non-primary RF signals may cause spourious contributions in the error signals,
including the other PDs.

ELOG V3.1.3-